### v2

# PAC-Bayesian Reinforcement Learning Trains Generalizable Policies

Abstract
We derive a novel PAC-Bayesian generalization bound for reinforcement learning that explicitly accounts for Markov dependencies in the data, through the chain’s mixing time. This contributes to overcoming challenges in obtaining generalization guarantees for reinforcement learning, where the sequential nature of data breaks the independence assumptions underlying classical bounds. Our bound provides non-vacuous certificates for modern off-policy algorithms like Soft Actor-Critic. We demonstrate the bound’s practical utility through PB-SAC, a novel algorithm that optimizes the bound during training to guide exploration. Experiments across continuous control tasks show that our approach provides meaningful confidence certificates while maintaining competitive performance.
1 Introduction
Deploying reinforcement learning (RL) algorithms in real environments and safety-critical applications requires high confidence that the learned policies will generalize beyond the training data. Although the statistical learning literature has proved rigorous generalization guarantees for supervised learning, and despite progress in generalization guarantees for deep learning models (Pérez-Ortiz et al., 2021) in this setting, some challenges prevent extending to other settings; particularly pertaining to data assumptions.
In RL, an outstanding challenge stems from the fact that algorithms learn from sequential, temporally correlated trajectories that break the standard assumption of independent and identically distributed (i.i.d.) data underlying most classical generalization theory. Since RL trajectories exhibit strong temporal dependencies, where future states depend on past actions and the evolving policy, classical analyses that rely on the i.i.d. assumption, and their associated sharp concentration bounds, cannot be directly applied to provide meaningful generalization guarantees in this setting.
Facing the aforementioned challenge, the PAC-Bayesian computations framework, which uses PAC-Bayes bounds (McAllester, 1999; Catoni, 2007; Germain et al., 2015; Alquier, 2024) as optimization objectives and for computing high confidence certificates (Pérez-Ortiz et al., 2021) that hold at distribution level, emerges as a promising solution: the analysis maintaining distributions over hypotheses rather than a single model can be extended to temporally dependent data through appropriate concentration inequalities.
Various approaches have tried to relax the independent data condition, to handle data with dependence structures. Martingale-based methods (Seldin et al., 2011; 2012) extend PAC-Bayesian analysis to sequential settings by constructing martingale sequences from the data (e.g., value function errors, Bellman residuals), then applying martingale concentration inequalities such as Azuma-Hoeffding’s (Azuma, 1967) and Freedman’s (Freedman, 1975). Although these approaches are mathematically elegant, RL problems do not naturally yield martingales. Since RL data possess Markov structure by design, concentration inequalities tailored specifically for Markov chains can provide a more natural and better suited approach to this domain.
In this work, we present a novel PAC-Bayesian bound for RL that explicitly accounts for Markov dependencies through the chain’s mixing time [^1] The smallest number of steps required for the chain’s distribution to become nearly indistinguishable from its stationary distribution, regardless of its initial state. . Our key technical contribution integrates a bounded-differences condition on the negative empirical return with McDiarmid-type concentration inequality for Markov chains (Paulin, 2018), yielding a bound with explicit constants and improved scaling that avoids the vacuity of previous approaches (cf. Section 2.4). We demonstrate that this new bound is not merely a theoretical curiosity by introducing PB-SAC, an actor-critic algorithm that makes practical use of the bound as a live, optimizable objective. The algorithm periodically computes the numerical value of the PAC-Bayesian bound and uses it to guide learning through posterior sampling, transforming the generalization guarantee from a passive post-training evaluation tool into an active component of a learning algorithm.
Thus, the work at hand proposes the following contributions. From a theoretical perspective, we develop a PAC-Bayesian bound for RL with explicit mixing-time dependence and improved scaling with realistic trajectory lengths and discount factors relative to prior work, representing a step toward RL certificates. Algorithmically, we introduce PB-SAC, the first practical algorithm that puts into action a non-vacuous PAC-Bayesian bound as an optimizable objective in modern deep RL, including key innovations for stable optimization. Empirically, we demonstrate that bound values remain informative across continuous-control tasks while maintaining competitive performance with state-of-the-art methods.
Our approach establishes the first practical PAC-Bayesian computations framework for optimization and certification in modern RL, bridging the gap between learning theory and algorithmic practice in sequential decision-making.
2 Preliminaries
In this section, we briefly recall the definitions of reinforcement learning and statistical learning theory concepts we rely on throughout the paper. The exposition in this section is intentionally concise—the goal is to fix notation and state the learning theory principles that underpin our results.
2.1 Reinforcement Learning
Reinforcement Learning (RL) studies how an agent learns to make decisions through sequential interaction with an environment. Formally, the environment is modeled as a (possibly unknown) Markov Decision Process (MDP) denoted $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathbb{P},R,\gamma)$; where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $\mathbb{P}(s^{\prime}\mid s,a)$ the transition kernel, $R(s,a)$ the reward function bounded in $0,R_{\max}]$, and $\gamma\in(0,1)$ the discount factor. At each time $t$, the agent observes a state $S_{t}\in\mathcal{S}$, chooses an action $A_{t}\in\mathcal{A}$ according to a policy $\pi(a|s)$, receives a reward $R_{t+1}=R(S_{t},A_{t})$, and transitions to a new state $S_{t+1}\sim\mathbb{P}(\cdot\mid S_{t},A_{t})$.
The learning agent’s objective is, at least in principle, to maximize the expected discounted return
$V_{\pi}(s)=\mathbb{E}_{\pi,\mathbb{P}}\!\bigl[G_{t}\mid S_{t}=s\bigr],$ (1)
where
$G_{t}=\sum_{k=0}^{\infty}\gamma^{k}\,R_{t+k+1}.$ (2)
The function $V_{\pi}(\cdot)$ is the state–value function. It is well known that the optimal value function $V^{\star}(s)=\sup_{\pi}V_{\pi}(s)$ satisfies the Bellman optimality equation
$V^{\star}(s)=\max_{a\in\mathcal{A}}\Bigl\{R(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim\mathbb{P}}\bigl[V^{\star}(s^{\prime})\mid s,a\bigr]\Bigr\}.$ (3)
RL algorithms either learn directly a policy (policy–gradient and actor–critic methods, see e.g. Haarnoja et al. (2018); Konda & Tsitsiklis (1999); Sutton et al. (1999)), or learn an action–value function $Q_{\pi}(s,a)$ in value–based methods such as Q‐learning and its deep variants Mnih et al. (2013). Model–free approaches do not use an explicit model of $\mathbb{P}$, while model–based methods leverage or learn a transition model to plan.
2.2 Probably Approximately Correct (PAC) Bounds
In a supervised learning task, the instance space is taken to be the product $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is the feature space and $\mathcal{Y}$ the label space; commonly $\mathcal{Y}\subseteq\mathbb{R}$ for regression, and $\mathcal{Y}\subseteq\mathbb{N}$ for classification problems. In this setting, a learning algorithm is a function that takes in a training sample $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{m}$ and returns a prediction function $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, also referred to as a hypothesis, parametrized by $\theta\in\Theta$, where $\Theta\subset\mathbb{R}^{p}$ denotes the set of all admissible parameter vectors. An unknown data distribution $\mathcal{D}$ over $\mathcal{Z}$ is postulated, with $\mathcal{D}_{\mathcal{X}}$ denoting the marginal distribution on $\mathcal{X}$, and training samples $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{m}$ are such that each pair $(\bm{x}_{i},y_{i})\in\mathcal{Z}$ is an independent and identically distributed (i.i.d.) random draw from $\mathcal{D}$, that is, $S\sim\mathcal{D}^{\otimes m}:=\mathcal{D}\otimes\cdots\otimes\mathcal{D}$ ($m$ copies). The “quality” of a hypothesis $f_{\theta}$ is typically assessed through a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, which quantifies the discrepancy between predicted and true outputs (labels). Two important functionals associated to a hypothesis are its true risk on the distribution $\mathcal{D}$, and its empirical risk on the sample $S$, which are respectively given by
$\mathcal{L}(\theta)$ $=\underset{{(x,y)\sim\mathcal{D}}}{\mathbb{E}}\bigl[\ell\bigl(f_{\theta}(\bm{x}),y\bigr)\bigr],$ (4)
$\hat{\mathcal{L}}_{S}(\theta)$ $=\frac{1}{m}\sum_{i=1}^{m}\ell\bigl(f_{\theta}(\bm{x}_{i}),y_{i}\bigr).$ (5)
In supervised machine learning, the goal is to learn a hypothesis $f_{\theta}$ that accurately predicts labels $y\in\mathcal{Y}$ for given inputs $\bm{x}\in\mathcal{X}$. Since training is based on a finite dataset, a central question is: how can we ensure that the learned function $f_{\theta}$ will perform well on unseen data?
Probably Approximately Correct (PAC) learning answers this question via probability inequalities saying that, under suitable restrictions, for a chosen level $\delta$ within $(0,1)$ we have
$\Pr_{S\sim\mathcal{D}^{m}}\bigl\{\mathcal{L}(\theta)\leq\hat{\mathcal{L}}_{S}(\theta)+\epsilon\bigr\}\geq 1-\delta.$
Choosing a small $\delta$ then translates into a high-confidence bound for $\mathcal{L}(\theta)$. Concrete PAC bounds specify how large $m$ must be (or how large the gap $\epsilon$ can be) in terms of properties of the hypothesis class, e.g. VC‑dimension, Rademacher complexity, stability, compression, etc.
All of the classical PAC bounds treated $f_{\theta}$ as a deterministic output of the algorithm.
2.3 PAC-Bayesian Bounds
[…]
$\underset{\theta\sim\rho}{\mathbb{E}}[\mathcal{L}(\theta)]$ $\leq\underset{\theta\sim\rho}{\mathbb{E}}[\hat{\mathcal{L}}_{S}(\theta)]+\frac{1}{\kappa}\Bigl(\mathrm{KL}(\rho\|\mu)+\ln\tfrac{1}{\delta}+\Psi_{\ell,\mu}(\kappa,m)\Bigr)$ (6)
with
$\Psi_{\ell,\mu}(\kappa,m)=\ln\underset{\theta\sim\mu}{\mathbb{E}}\;\underset{S\sim\mathcal{D}^{\otimes m}}{\mathbb{E}}\Bigl[\exp\bigl(\kappa\,\bigl(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{S}(\theta\bigr)\bigr)\Bigr].$
Compared with classical PAC guarantees, PAC‑Bayes offers two advantages that are critical for reinforcement learning: (1) Data‑dependent priors (Parrado-Hernández et al., 2012) —when $\mu$ can itself depend on previous data, e.g. earlier tasks or behavioural trajectories (Zhang et al., 2025), the bound adapts to the knowledge already acquired, tightening $\mathrm{KL}(\rho\|\mu)$; and (2) Fine‑grained control via $\Psi$ —by tailoring the concentration inequality used to upper‑bound $\Psi$ one can incorporate dependence structures such as martingales (Seldin et al., 2012; 2011), $\beta$‑mixing (Ralaivola et al., 2009; Abélès et al., 2025) sequences or Markov chains (Fard et al., 2011; Tasdighi et al., 2025). The latter is exactly the scenario in which RL trajectories are collected.
2.4 Previous PAC-Bayesian Bounds in RL
Early efforts to apply PAC-Bayesian theory to RL, notably the works of Fard & Pineau (2010) and Fard et al. (2011), established the framework’s viability for model selection in batch settings. However, their bounds suffered from poor scaling with the discount factor, rendering them numerically vacuous for problems with long effective horizons typical in modern RL (Tasdighi et al., 2025). Contemporary approaches have repurposed PAC-Bayes bounds for other algorithmic goals. Recent works have used them to derive training objectives for deep exploration (Tasdighi et al., 2025) or as regularizers for lifelong learning (Zhang et al., 2025). While these works demonstrate the versatility of PAC-Bayesian computations for algorithm design, they sidestep the original goal of providing tight, computable certificates for modern deep RL agents. This reveals a fundamental gap: although PAC-Bayesian theory holds promise for certified RL, existing approaches either yield vacuous bounds or repurpose them for different objectives. To the best of our knowledge, no practical algorithm has leveraged non-vacuous PAC-Bayesian bounds as live performance certificates within modern RL frameworks to date.
3 A new PAC-Bayesian Bound for RL
We now present our main theoretical contribution: a new PAC-Bayes generalization bound for RL that explicitly accounts for temporal dependencies through the mixing time of the underlying Markov chain.
3.1 Problem Setup
As outlined earlier, our objective is to establish a high‑probability PAC‑Bayes value‑error bound for a policy operating in a Markov decision process (MDP) when the training data are dependent trajectories—possibly gathered under an off‑policy algorithm. In this section, we begin by fixing notation and then present the main results; all proofs are deferred to Appendix B.
Let $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathbb{P},R,\gamma)$ be a discounted MDP, where $\mathcal{S}$ and $\mathcal{A}$ are the state and action spaces, $\mathbb{P}$ is the transition kernel, $R$ is the reward function such that $R_{t}\in[0,R_{\max}]$, and $\gamma\in(0,1)$ is the discount factor. A policy $\pi_{\theta}$ induces a (not necessarily time-homogeneous) Markov chain $\xi=(S_{1},A_{1},R_{1},S_{2},\dots,S_{H})$, where $H\leq\infty$ is the trajectory horizon (finite or infinite). Our analysis naturally extends to the infinite-horizon case ($H=\infty$). The distribution of this chain is determined by the initial state distribution $\nu$, transition probabilities $\mathbb{P}$, stochastic policy $\pi_{\theta}$, and random rewards $R$ as described next.
We assume access to a dataset $\mathfrak{D}=\{\xi^{(1)},\dots,\xi^{(T)}\}$ of $T$ trajectories (i.e., $N=HT$ transitions in total), collected using a behavior policy $\pi_{b}$. An evaluation policy $\pi_{\theta}$ is parameterized by $\theta\in\Theta$, the parameters $\theta$ are drawn from a distribution $\rho\in\mathcal{P}(\Theta)$, where $\Pi=\{\pi_{\theta}:\theta\in\Theta\}$ denotes the policy class. Henceforth, we write $\xi\sim\mathcal{M}$ (resp. $\mathfrak{D}\sim\mathcal{M}^{(T)}$) to denote sampling a trajectory (resp. a set $\mathfrak{D}$ of $T$ trajectories) under the environment dynamics $\mathbb{P}$, initial state distribution $\nu$, policy $\pi_{b}$, and reward function $R$, in order to avoid notational overload.
We define the discounted return of a trajectory $\xi$ and the value of policy $\pi_{\theta}$ respectively as:
$G(\xi)=\sum_{k=0}^{H-1}\gamma^{k}R_{k+1}\quad\text{and}\quad V_{\pi_{\theta}}=\mathbb{E}_{\xi\sim\mathcal{M}}[G(\xi)]$ (7)
We now define the expected (true) loss and its empirical counterpart using importance sampling to account for the off-policy data distribution:
$\mathcal{L}(\theta)=-\underset{\xi\sim\mathcal{M}}{\mathbb{E}}\left[\frac{\pi_{\theta}(\xi)}{\pi_{b}(\xi)}G(\xi)\right]=-\underset{\mathfrak{D}\sim\mathcal{M}^{(T)}}{\mathbb{E}}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)],$ (8)
$\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)=-\frac{1}{T}\sum_{j=1}^{T}w_{j}(\theta)G(\xi^{(j)}),$ (9)
where $w_{j}(\theta)=\frac{\pi_{\theta}(\xi^{(j)})}{\pi_{b}(\xi^{(j)})}$ is the likelihood ratio of trajectory $\xi^{(j)}\sim\pi_{b}$. To ensure theoretical stability, we assume the weighted rewards are bounded such that the weighted return remains within the range consistent with $R_{\max}$ (e.g., via clipping large importance weights, a standard practice in off-policy evaluation).
Following PAC-Bayesian folklore, we endow the parameter space $\Theta$ with a distribution $\mu\in\mathcal{P}(\Theta)$ playing the role of a prior selected independently of data (or partial dependence on it, cf. Parrado-Hernández et al. (2012); Pérez-Ortiz et al. (2021)), and a posterior $\rho\in\mathcal{P}(\Theta)$ chosen after observing $\mathfrak{D}$. This formalism enables reasoning about randomized policies drawn from $\rho$ with guarantees based on their divergence from $\mu$. Crucially for our analysis, changing one transition in the data results in a quantifiable bounded effect on the empirical loss defined in equation 8:
Lemma 3.1 (Bounded differences).
Let $\mathfrak{D}$ be a set of trajectories and $\theta\in\Theta$ be fixed policy parameters. Suppose we form $\bar{\mathfrak{D}}$ by changing one transition, say the transition at time step $h\in[H]$ of trajectory $j\in[T]$, where $\xi^{(j)}_{h}=(s,a,r,s^{\prime})$ is replaced with $\bar{\xi}^{(j)}_{h}=(\bar{s},\bar{a},\bar{r},\bar{s}^{\prime})$. Then, there exists $c\in\mathbb{R}_{+}^{H\times T}$ such that
$\hskip-5.0pt\bigl|\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\hat{\mathcal{L}}_{\bar{\mathfrak{D}}}(\theta)\bigr|\leq\sum_{h^{\prime}=1}^{H}\sum_{j^{\prime}=1}^{T}c_{(h^{\prime},j^{\prime})}\,\mathbb{I}\Bigl[\xi^{(j^{\prime})}_{h^{\prime}}\neq\bar{\xi}^{(j^{\prime})}_{h^{\prime}}\Bigr$] (10)
Intuitively, $c_{(h,j)}$ quantifies the transition‑level influence of altering the (h,j)(h,j)‑th state–action–reward tuple on the average return. A complete derivation—including a justification of why this bound covers propagation of the perturbed transition to future steps—is given in Appendix B.2. The result yields the explicit vector $c$ used in Theorem 3.2, namely
$c_{(h,j)}=\frac{\gamma^{h-1}R_{\max}}{T},$ (11)
and its norm
$\|c\|^{2}=\frac{R_{\max}^{2}}{T(1-\gamma^{2})}\bigl(1-\gamma^{2H}\bigr).$ (12)
3.2 Main Result
The above bounded-differences property (Lemma 3.1) is precisely what allows us to apply concentration inequalities to dependent data. Importantly, the proof strategy is leveraging Paulin (2018)’s extension of McDiarmid’s inequality to Markov chains, which provides concentration for functions satisfying bounded differences on Markovian sequences. The key insight is that while the transitions are temporally dependent, the bounded-differences condition with explicit constants $\|c\|^{2}$ enables us to control how perturbations propagate through the dependency structure.
Applying Paulin (2018)’s concentration result yields a tail bound on the deviation $\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)$ (see inequality equation 20) that depends explicitly on the mixing time $\tau_{\min}$ of the policy-induced Markov chain. $\tau_{\min}$ is the smallest number of steps after which the distribution of the chain’s state is, in a statistical sense, nearly indistinguishable from its long‑run or stationary distribution in total variation distance, no matter where the chain started. In other words, it measures how quickly the chain “forgets” its initial state and becomes well mixed. Combining the tail bound from inequality equation 20 with the standard PAC-Bayesian change-of-measure technique gives our main result:
Theorem 3.2.
Let the weighted reward function be bounded in $0,R_{max}]$ and let $\mathcal{M}$ be a (not necessarily time-homogeneous) Markov Decision Process (MDP) induced by any policy $\pi_{b}$ such that it satisfies $\tau_{\min}<+\infty$. For any prior $\mu$ over $\Theta$, any posterior $\rho$ chosen after interacting with the environment, and any $\delta\in(0,1)$, with probability at least $1-\delta$ over the sample $\mathfrak{D}$ of $T$ trajectories with time horizon $H$, :
$\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)]$ $\leq\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]+\sqrt{\frac{R_{\max}^{2}(1-\gamma^{2H})}{T(1-\gamma^{2})}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}.$ (13)
The bound in Theorem 3.2 can be straightforwardly converted to a PAC-Bayes lower bound on the true expected value function $\mathbb{E}_{\theta\sim\rho}\left[V_{\pi_{\theta}}\right]$ (see Appendix B, inequality equation 31), using the fact that $\mathcal{L}(\theta)=-V_{\pi_{\theta}}$ (equation 8, equation 9) and a simple rearrangement of terms. The true expected value is lower-bounded by an empirical estimate minus an uncertainty term that accounts for limited data ($1/T$), temporal correlations ($\tau_{\min}$), and posterior complexity ($\mathrm{KL}(\rho\|\mu)$). This interpretation suggests a natural approach to policy optimization: select the posterior $\rho$ that maximizes this lower bound (equivalently, minimizes the upper bound in inequality equation 13). Such a strategy would automatically balance exploitation (maximizing the empirical value) and theoretically-justified exploration (accounting for uncertainty).
3.3 Key Improvements and Discussion
Improved Horizon Dependence.
Previous PAC-Bayes bounds for RL suffer from prohibitive scaling with the effective horizon $1/(1-\gamma)$. Fard et al. (2011) bound the value error via the Bellman error, a conversion (their Eq. 7) that degrades sample complexity to $\mathcal{O}((1-\gamma)^{-4})$. Tasdighi et al. (2025) face comparable constraints. For $\gamma=0.99$, such scaling renders these bounds practically vacuous. In contrast, our transition-level analysis bounds the value error directly via the concentration of discounted returns, achieving a significantly tighter scaling of $\mathcal{O}((1-\gamma)^{-1})$. This improvement arises because our sensitivity term scales with $\sum^{H}\gamma^{2h}\approx(1-\gamma)^{-1}$ (Appendix B.2), avoiding the $(1-\gamma)^{-2}$ penalty inherent to the two-step derivation from the Bellman error. Specifically, we require only $T\gtrsim\frac{R_{\max}^{2}\tau_{\min}}{1-\gamma^{2}}$ trajectories.
Explicit Mixing Time Dependence.
Unlike bounds for general mixing processes that depend on abstract coefficients, our result features explicit dependence on $\tau_{\min}$, the mixing time of the policy-induced Markov chain. This quantity is well studied and has a clear interpretation in terms of environment dynamics. While extending concentration inequalities from independent to mixing settings is sometimes viewed as straightforward, the reality involves several subtle challenges. Although the distribution of $X_{t}$ becomes close to the stationary distribution $\pi$ after $\tau_{\min}$ steps, this does not guarantee that $X_{t}$ is approximately independent of $X_{0}$, let alone that consecutive states $X_{t}$ and $X_{t+1}$ are independent. The dependence between observations decays exponentially with the mixing time, but achieving approximate independence typically requires waiting several multiples of $\tau_{\min}$, not just $\tau_{\min}$ itself.
Moreover, applying McDiarmid-type inequalities to RL requires establishing that the negative empirical return satisfies a bounded-differences condition with explicit, tractable constants (Lemma 3.1). This necessitates careful analysis of how perturbations at individual transitions propagate through the sequential structure to affect future states and rewards. Our transition-level analysis directly addresses this challenge by quantifying the error propagation through the Markov dependency structure. The proof is provided in Appendix B.4.
Practical Tractability and Robustness.
The bound requires estimating $\tau_{\min}$, which presents both computational and robustness considerations. We estimate mixing time using autocorrelation decay of the reward signal, as this can be computed from streaming trajectories without storing full visitation counts. Alternative approaches utilizing the pseudo-spectral gap Karagulyan & Alquier (2025) can offer tighter bounds when the gap is estimable. However, constructing fully empirical estimators for these gaps currently requires finite state spaces or specific parametric assumptions (e.g., AR(1)), rendering them intractable for general deep reinforcement learning in continuous environments.
Our estimation approach is robust to errors in a specific direction: if we overestimate $\tau_{\min}$, our bound remains valid but becomes looser, which is harmless in practice. However, underestimation can be problematic as it leads to overconfidence. To mitigate this, we compute autocorrelation from additional sources (e.g., state features) to cross-validate mixing time estimates. In practice, we err on the side of caution by using a conservative initial estimate and taking the maximum with the latest autocorrelation estimation, trading some tightness for reliability. Empirical analysis can be found in Section 5.4.
4 PB-SAC: A Practical Algorithm for PAC-Bayesian Reinforcement Learning
Translating the theoretical PAC-Bayes bound into a practical learning algorithm requires addressing challenges of maintaining posterior distributions over policy parameters in deep RL. The novel algorithm proposed here, PAC-Bayes Soft Actor-Critic (PB-SAC), builds upon SAC while integrating PAC-Bayesian bounds and posterior-guided exploration (pseudo-code and illustrative figure in Appendix F).
PB-SAC’s central insight is that policy parameters always represent the posterior mean, updated through standard SAC gradients during regular training, ensuring most learning follows proven SAC dynamics while periodic PAC-Bayesian updates refine the posterior and guide exploration. Following Zhang et al. (2025), we maintain a diagonal Gaussian posterior $\rho(\theta)=\mathcal{N}(\upsilon,\operatorname{diag}(\sigma^{2}))$ over flattened policy parameters with learnable mean $\upsilon$ and standard deviation $\sigma$. The prior $\mu$ undergoes periodic moving average updates toward the current posterior with linear decay, preventing KL divergence explosion while preserving bound validity and maintaining exploration capability as the prior stabilizes during training.
4.1 Posterior-Guided Exploration
During exploration, PB-SAC leverages the posterior distribution to implement uncertainty-driven exploration. Rather than standard $\epsilon$-greedy exploration, the algorithm samples policies from the posterior and selects actions that maximize Q-values under posterior uncertainty:
$\theta_{explore}\leftarrow\arg\max_{\theta_{i}\in\mathfrak{T}}Q(s,\pi_{\theta}(s))$
such that $\mathfrak{T}$ is a set of policy parameter vectors drawn from $\rho$. This posterior-guided exploration naturally balances exploitation of the mean policy (the current actor) when it yields the highest value, with exploration of alternative policies in regions of high posterior uncertainty where potentially superior policies may exist. This approach provides theoretical grounding for the exploration strategy through PAC-Bayesian computations, ensuring that exploration is guided by uncertainty quantification rather than arbitrary randomness.
4.2 Alternating Optimization via PAC-Bayes-$\lambda$
The core challenge in optimizing our PAC-Bayesian bound (Theorem 3.2) lies in its structure: while the KL divergence term $\mathrm{KL}(\rho\|\mu)$ is convex in its first argument (the posterior parameters), the square root composition in the bound is not guaranteed to be convex, potentially leading to optimization difficulties. To address this challenge, we use variational approximation. By applying the identity $\sqrt{x}=\inf_{\lambda>0}(\frac{x}{2\lambda}+\frac{\lambda}{2})$, we recover the convex objective $\mathcal{J}(\rho,\lambda)$ in equation 14. This objective has a structure remarkably similar to the PAC-Bayes-$\lambda$ bound of Thiemann et al. (2017), derived from the classical PAC-Bayes-kl bound for majority vote learning (see, e.g., Germain et al. (2015), Corollary 21; Seeger (2003); Langford (2005)). Both approaches share the key insight of introducing an auxiliary trade-off parameter $\lambda$ that transforms non-convex objectives into quasi-convex relaxations amenable to stable alternating optimization.
Our implementation optimizes the following PAC-Bayes-$\lambda$ objective to obtain the posterior $\rho$ then substitutes this optimized posterior into inequality equation 13. As shown in our proof B.5.4, the main bound achieves its minimum over $\lambda^{*}$, yielding a tighter certificate:
$\mathcal{J}(\rho,\lambda)=\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]+\frac{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}{2\lambda}+\frac{\lambda}{2}$ (14)
This decomposition enables stable alternating optimization: we optimize posterior parameters for fixed $\lambda$, then compute optimal closed form $\lambda$ for fixed posterior parameters (equation equation 35), ensuring convergence and maintaining the theoretical guarantees of our PAC-Bayesian bound.
4.3 Policy-level REINFORCE trick
Even with the above decomposition, we cannot straightforwardly compute the gradient
$\nabla_{(\upsilon,\sigma)}\mathbb{E}_{\theta\sim\rho}\left[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)\right]$
because sampling cannot occur inside the gradient operation. To address this challenge, we employ a two-stage approach. First, we collect fresh rollouts during the PAC-Bayes update cycle using the mean policy, then sample policies from $\rho$ and evaluate their discounted returns on these rollouts using importance sampling. This ensures that bound computation accounts for distributional shift between the data-generating policy and the current posterior distribution. With estimated returns for each sampled policy, we apply the log-likelihood trick (REINFORCE, Williams (1992)) at the policy level rather than the traditional action level. We prove this extension in Appendix E. This technique allows us to exchange gradient and expectation operations:
$\mathbb{E}_{\theta\sim\rho}\left[\nabla_{(\upsilon,\sigma)}\log\mathbb{P}_{\upsilon,\sigma}(\theta)\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)\right],$
yielding a tractable sampling-based gradient estimator.
4.4 Adaptive Sampling
Our most critical innovation addresses the actor-critic misalignment problem that arises after PAC-Bayesian updates. When the posterior mean shifts significantly, critics become misaligned with the new policy distribution, creating instability in training. Without this mechanism, ablation studies (Figure 7(a)) reveal a characteristic sawtooth pattern: performance drops sharply immediately after each PAC-Bayesian update, then gradually recovers until the next update cycle.
PB-SAC resolves this through adaptive sampling: immediately following PAC-Bayesian updates, we freeze the actor and employ high-rate posterior sampling (256 samples) to expose critics to the full posterior distribution for recalibration, then resume efficient learning with minimal sampling (1 sample, the posterior mean), achieving computational efficiency without sacrificing stability.
5 Experiments
We now turn to empirical validation across representative continuous control tasks to demonstrate that PAC-Bayesian computations for RL deliver on this promise while maintaining competitive learning performance.
5.1 Experimental Design
We evaluate PB-SAC on four MuJoCo continuous control environments (Towers et al., 2024; Tassa et al., 2018) spanning different complexity levels: HalfCheetah, Ant, Hopper, and Walker2d. Our evaluation tracks two critical metrics: PAC-Bayesian certificate evolution (Figure 1, right) and learning performance relative to baselines (Figure 1, left).
We run PB-SAC with PAC-Bayesian updates every 20,000 steps, employing our adaptive sampling (512 posterior samples during critic adaptation, posterior mean alone during regular training). We compare against vanilla SAC using identical network architectures and include PBAC Tasdighi et al. (2025), a PAC-Bayes deep exploration method, despite its primary strengths manifesting in sparse reward settings. We justify this comparison by evaluating our algorithm in the same sparse-reward setting, with results reported in Appendix G. Additional ablation studies examining the impact of our algorithmic components (adaptive sampling, posterior-guided exploration, and mixing time estimation) are also provided in Appendix G. Table 1 summarizes all hyperparameters.
(a) Algorithm comparisons
(b) PAC-Bayes analysis
Figure 1: (a) Performance comparison between our   PB-SAC, its baseline   SAC, and   PBAC from Tasdighi et al. (2025); (b) PAC-Bayes analysis of PB-SAC across environments. The empirical discounted return (dashed line) corresponds to $\mathbb{E}_{\theta\sim\rho}[-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]$, and the certified discounted return (solid line) corresponds to the lower bound on $\mathbb{E}_{\theta\sim\rho}[-\mathcal{L}(\theta)]$ provided by Theorem 3.2 (after rearranging the terms).
5.2 The Tightening of Performance Certificates
Figure 1(b) reveals the most compelling aspect of our results: the PAC-Bayesian bounds consistently tighten throughout training, tracking the improvement in learned policies. In HalfCheetah, the bounds become informative within 100k steps and continue tightening as performance improves, demonstrating that our certificates provide genuine confidence estimates for high-performing policies rather than vacuous guarantees. Particularly noteworthy is the bounds’ behavior during performance fluctuations: they appropriately widen during periods of high variance whilst tightening when performance stabilizes (e.g., HalfCheetah between 100k and 200k steps). This stability stems from our decaying moving average prior updates, which prevent KL explosion in early stages; without this mechanism, KL explosion hinders learning by pulling the posterior back towards the prior rather than allowing improvement.
Hopper, Walker2d (Figure 4), and Ant exhibit similar bound evolution, with certificates tracking performance and reflecting meaningful task completion. Notably, Ant—the most challenging test case due to its 3D dynamics and large state space—demonstrates that our bounds remain meaningful even for complex, high-dimensional continuous control tasks. The bounds capture the qualitative behavior predicted by our theory.
5.3 Empirical Validation Across Environments
The learning curves in Figure 1(a) address a fundamental concern about PAC-Bayesian RL: whether theoretical guarantees necessitate performance sacrifices. Our results demonstrate that PB-SAC maintains competitive performance with SAC across all tested environments while providing formal certificates. In HalfCheetah and Ant, PB-SAC matches or slightly exceeds SAC performance. In Hopper and Walker2d, both algorithms achieve strong performance, with PB-SAC maintaining stability throughout training. This validates our architectural choices, particularly the policy-posterior synchronization that ensures most learning follows proven SAC dynamics, with periodic PAC-Bayesian updates serving as refinements rather than disruptions, guided by the adaptive sampling that prevents critic destabilization.
PBAC, designed specifically for deep exploration through multi-objective optimization (diversity, coherence, and propagation), underperforms both methods on dense-reward tasks. To ensure a fair comparison—since PBAC’s primary strengths manifest in sparse-reward settings—we evaluated PB-SAC on the sparse-reward task Ant (very delayed) from PBAC’s original work (Tasdighi et al., 2025). Results (Appendix G.1) demonstrate that PB-SAC matches and ultimately surpasses PBAC while outperforming vanilla SAC, confirming that our posterior-guided exploration provides meaningful benefits in sparse-reward settings without requiring PBAC’s high computational complexity (ensemble of critic networks). These results establish that our framework provides non-vacuous certificates and competitive performance across both dense and sparse reward tasks.
5.4 Mixing Time Analysis
A critical component of our framework is mixing time estimation, as underestimation of $\tau_{\min}$ can lead to overconfident bounds. To assess robustness, we conducted experiments with fixed mixing time estimates ranging from 1 (fastest mixing, potential underestimation) to 1000 (conservative overestimation for MuJoCo tasks) without dynamic updates during training. Figure 6 illustrates these results on HalfCheetah. As expected, smaller mixing time values produce tighter bounds [^2] Using a mixing time equal to 1 is equivalent to an i.i.d. bound, more specifically to one that can be derived from the original McDiarmid’s inequality (McDiarmid, 1989) (not the one we used). This demonstrates how tight such bounds can be., while larger values yield more conservative certificates. However, a notable finding is that overestimation proves less problematic than anticipated: while the bound remains theoretically valid but looser, the algorithm effectively optimizes it, maintaining meaningful certificates even with substantial overestimation. This robustness suggests that erring on the side of caution with conservative mixing time estimates is a viable practical strategy. In practice, we mitigate underestimation risk through cross-validation using autocorrelation from multiple signals (rewards and state features) and maintain monotonically increasing estimates by taking the maximum across updates. These results demonstrate that our framework provides reliable certificates across a wide range of mixing time specifications.
6 Conclusion
We introduced PB-SAC, a PAC-Bayesian soft actor-critic algorithm that bridges the gap between learning theory and practical deep RL by making practical use of non-vacuous performance certificates as active components of the learning process. Our theoretical contribution (a PAC-Bayes bound with explicit mixing-time dependence and improved discount factor scaling) enables meaningful generalization guarantees in sequential decision-making. Algorithmically, our key innovations include policy-posterior synchronization, adaptive sampling for critic adaptation, posterior-guided exploration, and prior moving average that maintains competitive performance while providing formal certificates.
Several limitations warrant consideration. Our approach requires accurate mixing time estimation (underestimation leads to overconfident bounds), faces computational overhead that may limit scalability to large networks, and uses KL divergence which, despite its analytical convenience, is unstable when distributions diverge significantly and does not respect the geometry of the parameter space (Viallard et al., 2023). Future work could explore efficient posterior approximations, meaningful representations (layer/parameter-level rather than full network posteriors), adaptive mixing time estimation, Wasserstein distance alternatives to KL (Amit et al., 2022; Haddouche & Guedj, 2023; Viallard et al., 2023), and extensions beyond actor-critic frameworks. Despite these constraints, our results believably establish PAC-Bayesian computations for RL as a viable approach for certified performance in modern deep reinforcement learning, particularly valuable in high-stakes applications requiring theoretical guarantees.
Acknowledgments
We gratefully acknowledge support from the CNRS/IN2P3 Computing Center (Lyon - France) for providing computing resources needed for this work.
Appendix A Mathematical Tools
Lemma A.1 (Markov’s Inequality).
For any random variable $X$ such that $\mathbb{E}[|X|]=\mu$, for any $a>0$, we have
$\mathbb{P}\{|X|\geq a\}\leq\frac{\mu}{a}.$
Lemma A.2 (Change of measure).
For any measurable function $f:\Theta\rightarrow\mathbb{R}$ and distributions $\mu,\rho\in\mathcal{P}(\Theta)$:
$\mathbb{E}_{\theta\sim\rho}[f(\theta)]\leq\mathrm{KL}(\rho\|\mu)+\ln\mathbb{E}_{\theta\sim\mu}[\exp(f(\theta))]$ (15)
where $\mathrm{KL}(\rho\|\mu)$ is the Kullback-Leibler divergence between $\rho$ and $\mu$.
A.1 Concentration for Markov chains via Marton coupling
We use Paulin’s extension of McDiarmid’s bounded‑difference inequality to Markov chains (Paulin, 2018). This extension provides concentration inequalities for functions of dependent random variables, with constants that depend on the mixing properties of the chain.
A.1.1 Marton coupling and mixing time
The key insight in Paulin’s approach (Paulin, 2018) is to use a coupling structure known as Marton coupling, which quantifies the dependency between random variables in a Markov chain. For a Markov chain $X=(X_{1},\ldots,X_{N})$ on state space $\Lambda=\Lambda_{1}\times\ldots\times\Lambda_{N}$, a Marton coupling provides a way to couple the distributions of future states conditioned on different past states.
Let $\tau(\varepsilon)$ denote the mixing time of the chain $X$ in total variation distance, defined as the minimal $t$ such that for every $1\leq i\leq N-t$ and $x,y\in\Lambda_{i}$:
$d_{TV}(\mathcal{L}(X_{i+t}|X_{i}=x),\mathcal{L}(X_{i+t}|X_{i}=y))\leq\varepsilon$ (16)
We define the normalized mixing time parameter $\tau_{\min}$ as:
$\tau_{\min}\;=\;\inf_{0\leq\varepsilon<1}\tau(\varepsilon)\Bigl(\tfrac{2-\varepsilon}{1-\varepsilon}\Bigr)^{2}$ (17)
A.1.2 McDiarmid’s inequality for Markov chains
For a function $f(X)$ satisfying the bounded-differences property: for any $x,y\in\Lambda$,
$f(x)-f(y)\leq\sum_{i=1}^{N}c_{i}\mathbb{I}[x_{i}\neq y_{i}]$ (18)
where $c\in\mathbb{R}^{N}_{+}$ and $\mathbb{I}[\text{condition}]$ is the indicator function, Paulin’s theorem gives:
$\Pr\bigl(|f(X)-\mathbb{E}f(X)|\geq t\bigr)\;\leq\;2\exp\bigl(-2t^{2}/\|c\|^{2}\tau_{\min}\bigr).$ (19)
The norm $\|c\|^{2}$ is defined as $\sum_{i=1}^{N}c_{i}^{2}$.
A.1.3 Application to bounded differences in MDPs
For Markov decision processes, this inequality is particularly useful when analyzing the difference between value functions. If perturbing a single transition can change the value by at most $c_{i}$, then the total effect on a function of trajectories is bounded by the above concentration inequality, with the mixing time of the MDP properly accounting for the propagation of the perturbation through future states.
Appendix B Derivation of PAC-Bayes Value-Error Bound for RL
B.1 Bounded-differences property for MDP trajectories
We begin by recalling the definitions of discounted return for a trajectory $\xi$ and the corresponding value function from Section 3:
$G(\xi)$ $=\sum_{k=0}^{H-1}\gamma^{k}R_{k+1}$
$V_{\pi_{\theta}}$ $=\mathbb{E}_{\xi\sim\mathcal{M}}[G(\xi)]$
As defined in equation 8 and equation 9, our empirical and expected losses are:
$\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)$ $=-\frac{1}{T}\sum_{j=1}^{T}G(\xi^{(j)})$
$\mathcal{L}(\theta)$ $=-\mathbb{E}_{\xi\sim\mathcal{M}}[G(\xi)]=-V_{\pi_{\theta}}$
To apply McDiarmid’s inequality for Markov chains, we must establish the bounded-differences condition for our empirical loss. Specifically, we need to show that replacing one transition in a trajectory affects $\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)$ by at most $\sum_{h=1}^{H}\sum_{j=1}^{T}c_{(h,j)}\mathbb{I}[\xi^{(j)}_{h}\neq\bar{\xi}^{(j)}_{h}$], where $c\in\mathbb{R}^{H\times T}_{+}$ and $\mathbb{I}$ is the indicator function.
B.2 Quantifying the impact of perturbed transitions
Suppose we replace a single transition at position $h$ in trajectory $j$. This perturbation alters the trajectory $\xi^{(j)}$ to $\bar{\xi}^{(j)}$ and effectively changes the cumulative importance weight associated with that trajectory.
By our assumption in Section 3.1, the weighted rewards are effectively bounded in $0,R_{\max}]$. Therefore, we can bound the sensitivity of the empirical loss by treating the weighted signal similarly to the on-policy case. The maximum change in the weighted discounted return due to a perturbation at step $h$ is bounded by the propagation of the error through the discount factor:
$|\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\hat{\mathcal{L}}_{\bar{\mathfrak{D}}}(\theta)|\leq\frac{\gamma^{h-1}R_{\max}}{T}=c_{(h,j)}$
B.3 Derivation of $\|c\|^{2}$ for the PAC-Bayes bound
To apply McDiarmid’s inequality for Markov chains as developed by Paulin (2018), we need to compute $\|c\|^{2}$:
$\lVert c\rVert^{2}$ $=\sum_{j=1}^{T}\sum_{h=1}^{H}c^{2}(h,j)$
$=\frac{R_{\max}^{2}}{T^{2}}\;\cdot T\sum_{h=1}^{H}\gamma^{\,2(h-1)}$
$=\frac{R_{\max}^{2}}{T}\;\underbrace{\sum_{h=0}^{H-1}\gamma^{\,2h}}_{\text{finite geometric series}}$
$=\frac{R_{\max}^{2}}{T}\cdot\frac{1-\gamma^{2H}}{1-\gamma^{2}}\,.$
For infinite-horizon settings where $H\rightarrow\infty$ and $\gamma<1$, the series converges to $1/(1-\gamma^{2})$, this simplifies to
$\lVert c\rVert^{2}=\frac{R_{\max}^{2}}{T(1-\gamma^{2})}.$
B.4 Full accounting of perturbation propagation effects
A critical question is whether our derivation of $\|c\|^{2}$ fully accounts for the propagation of perturbations through the trajectory. Since a perturbation at step $h$ in trajectory $j$ affects all subsequent transitions in that trajectory, the bounded-differences indicator is 1 for every $(h^{\prime},j)$ with $h^{\prime}\geq h$.
For a perturbation at step $h$ in trajectory $j$, the sum of corresponding coefficients is:
$\sum_{h^{\prime}=h}^{H}c_{(h^{\prime},j)}=\frac{R_{\max}}{T}\sum_{k=0}^{H-h}\gamma^{h-1+k}=\frac{R_{\max}\gamma^{h-1}}{T}\cdot\frac{1-\gamma^{H-h+1}}{1-\gamma}$
The actual maximum change in discounted return from this perturbation (worst case: reward changes from 0 to $R_{\max}$) is:
$|G(\xi^{(j)})-G(\bar{\xi}^{(j)})|\leq R_{\max}\gamma^{h-1}\sum_{k=0}^{H-h}\gamma^{k}=R_{\max}\gamma^{h-1}\cdot\frac{1-\gamma^{H-h+1}}{1-\gamma}$
When divided by $T$ (because $\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)$ averages over $T$ trajectories), we get exactly the same quantity as the sum of coefficients above. Therefore, the bounded-differences condition holds with equality, confirming that our derivation of $\|c\|^{2}$ fully accounts for all propagation effects.
This careful accounting of propagation effects allows us to apply McDiarmid’s inequality for Markov chains to obtain the PAC-Bayes bound in Theorem 3.2 with the correct constants.
B.5 Derivation of the PAC-Bayes Bound
Having established the bounded-differences property and quantified the impact of perturbations via $\|c\|^{2}$, we now derive the PAC-Bayes bound on the expected difference between empirical and true losses.
B.5.1 From McDiarmid to moment generating function
McDiarmid’s inequality for Markov chains (inequality 19) provides a concentration inequality on the deviation between empirical and expected losses. From this, we can derive a bound on the moment generating function (MGF) as shown by Paulin (2018):
Lemma B.1 (MGF bound for Markov chains).
For any $\kappa>0$ and policy parameters $\theta\in\Theta$:
$\mathbb{E}_{\mathfrak{D}\sim\mathcal{M}^{(T)}}\left[\exp\left(\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))\right)\right]\leq\exp\left(\frac{\kappa^{2}\|c\|^{2}\tau_{\min}}{8}\right)$ (20)
where $\tau_{\min}$ is the mixing time of the Markov chain induced by a behavior policy $\pi_{b}$.
B.5.2 Final bound via Sub-Gaussian Quadratic Moment
Instead of optimizing the free parameter $\kappa$ in equation 42, which would require a data-independent grid search and a union bound (Alquier, 2024), we leverage the sub-Gaussian nature of our generalization error to derive a tighter bound directly.
From Lemma B.1 (Equation 20), we established that the generalization error $\Delta(\theta)=\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)$ satisfies the sub-Gaussian property with variance proxy $\sigma_{proxy}^{2}$:
$\mathbb{E}_{\mathfrak{D}}\left[\exp(\kappa\Delta(\theta))\right]\leq\exp\left(\frac{\kappa^{2}\|c\|^{2}\tau_{\min}}{8}\right)=\exp\left(\frac{\kappa^{2}\sigma_{proxy}^{2}}{2}\right),\quad\text{where }\sigma_{proxy}^{2}=\frac{\|c\|^{2}\tau_{\min}}{4}.$ (21)
We utilize the property that centered sub-Gaussian random variables satisfy a bound on their quadratic exponential moment. Specifically, following Wainwright (2019) (Theorem 2.6.IV) and the correction by Hellström & Durisi (2020; 2021), for a $\nu$-sub-Gaussian variable $X$, and for any $\lambda\in[0,1)$, we have $\mathbb{E}[\exp(\frac{\lambda X^{2}}{2\sigma_{proxy}^{2}})]\leq\frac{1}{\sqrt{1-\lambda}}$. Choosing $\lambda=\frac{1}{2}$ for simplicity:
$\mathbb{E}_{\mathfrak{D}}\left[\exp\left(\frac{\Delta(\theta)^{2}}{4\sigma_{proxy}^{2}}\right)\right]\leq\sqrt{2}.$ (22)
We now apply the PAC-Bayesian change of measure (Lemma A.2) to the function $f(\theta)=\frac{\Delta(\theta)^{2}}{4\sigma_{proxy}^{2}}$:
$\mathbb{E}_{\theta\sim\rho}\left[\frac{(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))^{2}}{4\sigma_{proxy}^{2}}\right]\leq\mathrm{KL}(\rho\|\mu)+\ln\mathbb{E}_{\theta\sim\mu}\left[\exp\left(\frac{\Delta(\theta)^{2}}{4\sigma_{proxy}^{2}}\right)\right].$ (23)
Taking the expectation with respect to the data distribution $\mathfrak{D}\sim\mathcal{M}^{(T)}$ and applying Jensen’s inequality to the concave function $\ln(\cdot)$ followed by Fubini’s theorem to swap expectations in the second term:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}\left[\frac{(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))^{2}}{4\sigma_{proxy}^{2}}\right]$ $\leq\mathbb{E}_{\mathfrak{D}}[\mathrm{KL}(\rho\|\mu)]+\ln\mathbb{E}_{\theta\sim\mu}\mathbb{E}_{\mathfrak{D}}\left[\exp\left(\frac{\Delta(\theta)^{2}}{4\sigma_{proxy}^{2}}\right)\right]$ (24)
[…]
Multiplying by $4\sigma_{proxy}^{2}$ and substituting $\sigma_{proxy}^{2}=\frac{\|c\|^{2}\tau_{\min}}{4}$:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}\left[(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))^{2}\right]$ $\leq\|c\|^{2}\tau_{\min}\left(\mathbb{E}_{\mathfrak{D}}[\mathrm{KL}(\rho\|\mu)]+\ln\sqrt{2}\right).$ (26)
Finally, we apply Jensen’s inequality to the left-hand side, using $(\mathbb{E}[X])^{2}\leq\mathbb{E}[X^{2}]$:
$\left(\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}\left[\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)\right]\right)^{2}\leq\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}\left[(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))^{2}\right].$ (27)
Taking the square root gives the final in-expectation bound:
$\left|\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]\right|\leq\sqrt{\|c\|^{2}\tau_{\min}\left(\mathbb{E}_{\mathfrak{D}}[\mathrm{KL}(\rho\|\mu)]+\frac{1}{2}\ln 2\right)}.$ (28)
B.5.3 High-probability bound
To obtain the high-probability result (Theorem 3.2), we apply Markov’s inequality directly to the random variable $Z=\exp\left(\frac{\Delta(\theta)^{2}}{4\sigma_{proxy}^{2}}\right)$ within the change of measure step, similar to Hellström & Durisi (2021, Eq. 3). With probability at least $1-\delta$:
$\mathbb{E}_{\theta\sim\rho}\left[(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))^{2}\right]\leq 4\sigma_{proxy}^{2}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right).$ (29)
Using Jensen’s inequality $(\mathbb{E}_{\theta\sim\rho}[\Delta(\theta)])^{2}\leq\mathbb{E}_{\theta\sim\rho}[\Delta(\theta)^{2}]$ and taking the square root:
$\left|\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]\right|\leq\sqrt{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}.$ (30)
Substituting $\|c\|^{2}$ from equation 12 yields the result in Theorem 3.2.
Recalling that $\mathcal{L}(\theta)=-V_{\pi_{\theta}}$ from equation 8, we obtain the PAC-Bayes lower bound on the expected value function:
$\mathbb{E}_{\theta\sim\rho}[V_{\pi_{\theta}}]\geq\mathbb{E}_{\theta\sim\rho}[-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]-\sqrt{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}.$ (31)
where he true expected value is lower-bounded by an empirical estimate minus an uncertainty term that accounts for limited data ($1/T$), temporal correlations ($\tau_{\min}$), and posterior complexity ($\mathrm{KL}(\rho\|\mu)$).
B.5.4 Optimization of the bound
The high-probability bound derived above takes the following form:
$\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)]\leq\underbrace{\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]}_{\text{Linear in }\rho}+\underbrace{\sqrt{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}}_{\text{Concave in KL (Non-convex in }\rho)}.$ (32)
Directly optimizing this objective with respect to the posterior $\rho$ is challenging because the square root of the KL divergence makes the overall objective non-convex. To address this, we employ a variational approximation to obtain a surrogate objective that is amenable to alternating convex optimization.
We utilize the algebraic identity for the square root function: for any $x\geq 0$, $\sqrt{x}=\inf_{\lambda>0}\left(\frac{x}{2\lambda}+\frac{\lambda}{2}\right)$. By identifying $x$ with the complexity term inside the square root, we obtain an upper bound on the true risk that depends on the auxiliary parameter $\lambda$:
$\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)]\leq\inf_{\lambda>0}\mathcal{J}(\rho,\lambda),$ (33)
where the surrogate objective $\mathcal{J}(\rho,\lambda)$ is defined as:
$\mathcal{J}(\rho,\lambda)=\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]+\frac{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}{2\lambda}+\frac{\lambda}{2}.$ (34)
This surrogate objective has two desirable properties that justify the alternating optimization procedure in our algorithm (Section 4):
Convexity in $\rho$: For a fixed $\lambda$, the term becomes a linear scaling of the KL divergence, which is convex in $\rho$. This allows for stable gradient-based updates of the posterior parameters.
Convexity in $\lambda$: For a fixed $\rho$, the function is strictly convex in $\lambda$, allowing for an analytic solution:
$\lambda^{*}=\sqrt{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}.$ (35)
Minimizing $\mathcal{J}(\rho,\lambda)$ via alternating updates is therefore mathematically equivalent to minimizing the rigorous high-probability bound derived in Section B.5.3.
Appendix C Alternative derivation via union bound and a grid over $\kappa$
Below, we provide another alternative derivation, keeping the parameter $\kappa$ and optimizing it via a grid search and union bound following the procedure from Alquier (2024).
C.0.1 PAC-Bayes change of measure
Starting from the MGF bound above (Lemma B.1), we can follow the standard PAC-Bayes derivation. Let $\Theta$ be our parameter space and let $\mu\in\mathcal{P}(\Theta)$ be a prior distribution over $\Theta$ chosen independently of the data. For any posterior distribution $\rho\in\mathcal{P}(\Theta)$ (which may depend on $\mathfrak{D}$), we apply the change of measure inequality that follows from Donsker–Varadhan variational formula (Donsker & Varadhan, 1975; 1983).
Let $f(\theta)=\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))$. Applying Lemma A.2:
$\mathbb{E}_{\theta\sim\rho}[\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))]$ $\leq\mathrm{KL}(\rho\|\mu)+\ln\mathbb{E}_{\theta\sim\mu}[\exp(\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)))]$ (36)
C.0.2 Combining with the MGF bound
Taking the expectation with respect to $\mathfrak{D}\sim\mathcal{M}^{(T)}$ on both sides:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))]$ $\leq\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]+\mathbb{E}_{\mathfrak{D}}\left[\ln\mathbb{E}_{\theta\sim\mu}[\exp(\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)))]\right]$ (37)
Applying Jensen’s inequality to the concave function $\ln(\cdot)$ yields:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))]$ $\leq\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]+\ln\mathbb{E}_{\mathfrak{D}}\left[\mathbb{E}_{\theta\sim\mu}[\exp(\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)))]\right]$ (38)
By Fubini-Tonelli’s theorem, we swap the order of expectations in the second term and apply the MGF bound from Lemma B.1:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))]$ $\leq\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]+\ln\mathbb{E}_{\theta\sim\mu}\mathbb{E}_{\mathfrak{D}}[\exp(\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)))]$ (39)
$\leq\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]+\ln\mathbb{E}_{\theta\sim\mu}\left[\exp\left(\frac{\kappa^{2}\|c\|^{2}\tau_{\min}}{8}\right)\right]$ (40)
$=\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]+\frac{\kappa^{2}\|c\|^{2}\tau_{\min}}{8}$ (41)
Dividing by $\kappa>0$:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]$ $\leq\frac{\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]}{\kappa}+\frac{\kappa\|c\|^{2}\tau_{\min}}{8}$ (42)
where The true expected value is lower-bounded by an empirical estimate minus an uncertainty term that accounts for limited data ($1/T$), temporal correlations ($\tau_{\min}$), and posterior complexity ($\mathrm{KL}(\rho\|\mu)$).
Appendix D A note on the Markov assumption for Bellman errors
The work of Tasdighi et al. (2025) makes an interesting theoretical contribution by modeling the sequence of Bellman errors as a Markov chain. While this approach provides valuable insights, it is worth examining the conditions under which this assumption holds.
We present a simple illustrative example that highlights when the Markov property may not apply to Bellman error sequences. Consider a basic MDP with four states $\{A,B,C,D\}$ and the following transition dynamics with discount factor $\gamma=0$:
  • State $A$ transitions to state $C$ with reward $r=0$
  • State $B$ transitions to state $D$ with reward $r=0$
  • State $C$ has a self-loop with reward $r=+1$
  • State $D$ has a self-loop with reward r=1r=-1
AABBCCDDr=0r=0r=0r=0r=+1r=+1r=1r=-1
Figure 2: A basic four-state MDP
Using a value function $V\equiv 0$ that assigns zero value to all states, we can compute the Bellman errors:
$\delta_{t}(A)$ $=r(A)+\gamma\max_{a}\mathbb{E}[V(s^{\prime})|s=A,a]-V(A)=0+0\cdot V(C)-0=0$ (43)
$\delta_{t}(B)$ $=r(B)+\gamma\max_{a}\mathbb{E}[V(s^{\prime})|s=B,a]-V(B)=0+0\cdot V(D)-0=0$ (44)
Both states $A$ and $B$ yield the same Bellman error $\delta_{t}=0$ at time $t$. However, the subsequent errors differ:
$\delta_{t+1}(C)$ $=r(C)+\gamma\max_{a}\mathbb{E}[V(s^{\prime})|s=C,a]-V(C)=+1+0\cdot V(C)-0=+1$ (45)
$\delta_{t+1}(D)$ $=r(D)+\gamma\max_{a}\mathbb{E}[V(s^{\prime})|s=D,a]-V(D)=-1+0\cdot V(D)-0=-1$ (46)
This example demonstrates that the current Bellman error value $\delta_{t}=0$ alone does not uniquely determine the distribution of $\delta_{t+1}$, which depends on the underlying state that generated the current error. Although this example used discount $\gamma=0$ for the sake of simplicity, it is worth noting that the same example can be done with $\gamma>0$, leading to the same conclusion.
This observation suggests that the Markov assumption for Bellman error sequences may require additional conditions or refinements to hold. Such considerations could be valuable for future theoretical developments building upon the framework proposed by Tasdighi et al. (2025).
Appendix E REINFORCE Trick for Policy-Level Gradients
We prove that the REINFORCE trick can be extended from action-level to policy-level gradients, enabling tractable optimization of expectations over policy parameters.
Theorem E.1 (Policy-Level REINFORCE).
For a continuous posterior distribution $\rho_{\upsilon,\sigma}(\theta)=\mathcal{N}(\upsilon,\operatorname{diag}(\sigma^{2}))$ over policy parameters $\theta$, the gradient of the expected return can be computed as:
$\nabla_{\upsilon,\sigma}\mathbb{E}_{\theta\sim\rho}\left[\mathbb{E}[R_{t}|\pi_{\theta}]\right]=\mathbb{E}_{\theta\sim\rho}\left[\nabla_{\upsilon,\sigma}\log\rho_{\upsilon,\sigma}(\theta)\cdot\mathbb{E}[R_{t}|\pi_{\theta}]\right]$ (47)
Proof.
Let $J(\theta)=\mathbb{E}[R_{t}|\pi_{\theta}]$ denote the expected return of a specific policy parameterization $\theta$. We express the objective as an expectation over the continuous posterior density $\rho_{\upsilon,\sigma}(\theta)$ and apply the log-derivative trick directly:
$\nabla_{\upsilon,\sigma}\mathbb{E}_{\theta\sim\rho}\left[J(\theta)\right]$ $=\nabla_{\upsilon,\sigma}\int_{\Theta}\rho_{\upsilon,\sigma}(\theta)J(\theta)d\theta$ (48)
$=\int_{\Theta}\nabla_{\upsilon,\sigma}\rho_{\upsilon,\sigma}(\theta)J(\theta)d\theta$ (49)
$=\int_{\Theta}\rho_{\upsilon,\sigma}(\theta)\frac{\nabla_{\upsilon,\sigma}\rho_{\upsilon,\sigma}(\theta)}{\rho_{\upsilon,\sigma}(\theta)}J(\theta)d\theta\quad\text{(Log-derivative identity)}$ (50)
$=\int_{\Theta}\rho_{\upsilon,\sigma}(\theta)\nabla_{\upsilon,\sigma}\log\rho_{\upsilon,\sigma}(\theta)J(\theta)d\theta$ (51)
$=\mathbb{E}_{\theta\sim\rho}\left[\nabla_{\upsilon,\sigma}\log\rho_{\upsilon,\sigma}(\theta)\cdot J(\theta)\right]$ (52)
Substituting $J(\theta)$ back into the expression yields the theorem statement. ∎
This result enables us to estimate the gradient via sampling: we sample policies $\{\theta_{i}\}$ from the posterior $\rho$, evaluate their expected returns, and compute the weighted gradient of the log-probability density. This extends the classical REINFORCE trick from action spaces to parameter spaces, allowing efficient optimization of posterior distributions over policies.
Appendix F PB-SAC Algorithm
This appendix presents the complete algorithmic details of our PAC-Bayes Soft Actor-Critic (PB-SAC) method. Figure 3 provides a high-level illustration of the algorithm flow described in Section 4, showing the interaction between the actor, critic, and bound computation components. The periodic posterior update (step 2) guides exploration through PAC-Bayes bounds (step 3) while maintaining standard SAC training (step 1) using an adaptive sampling (step 4). Section F.1 provides the complete pseudocode with implementation details, including the posterior-guided exploration mechanism, the alternating bound optimization, and the integration with the critic adaptation loop.
Figure 3: Illustration of our algorithm PB-SAC
F.1 Pseudo Code
Result: Policy $\pi_{\theta}$, posterior $\rho$, PAC-Bayes bound
1
Init: Actor $\pi_{\theta}$, Critics $Q_{1},Q_{2}$, replay buffer $\mathfrak{D}$
Init: Posterior $\rho(\theta)=\mathcal{N}(\upsilon,\operatorname{diag}(\sigma^{2}))$ where $\upsilon$ are the initial actor parameters
Init: Prior $\mu(\theta)=\mathcal{N}(\upsilon_{0},\operatorname{diag}(\sigma_{0}^{2}))$
Init: $\text{actor\_frozen}=\text{False}$, prior moving average decay $\iota=0.99$, posterior sampling rate $\epsilon_{\text{explore}}=0.9$ (linearly decayed)
2
3for $t=1,2,\ldots$ do
4 if not actor_frozen then
    /* Standard SAC + Posterior-Guided Exploration */
5    if $\text{random}()<\epsilon_{\text{explore}}$ then
       /* Select policy maximizing Q-value from posterior */
6         Sample multiple policies $\{\theta_{i}\}\sim\rho$
7
8       $\theta_{\text{explore}}\leftarrow\arg\max_{\theta_{i}}Q(s,\pi_{\theta}(s))$
9
10       $a\leftarrow\pi_{\theta_{\text{explore}}}(s)$
11
12      end if
13    else
       $a\leftarrow\pi_{\theta}(s)$
       // Current policy (posterior mean)
14
15      end if
16     Execute action $a$ and store transition in $\mathfrak{D}$
17
18    if $|\mathfrak{D}|\geq\text{batch\_size}$ then
19         Update critics $Q_{1},Q_{2}$ with standard SAC loss
20         Update actor $\pi_{\theta}$ with standard SAC loss
       $\upsilon\leftarrow\theta$
       // Sync posterior mean to current actor
21
22
23      end if
24
25   end if
 /* Critic Adaptation Phase (post PAC-Bayes update) */
26   Sample multiple policies $\{\theta_{i}\}\sim\rho$ with high sampling rate
27
28  Compute critic targets averaged over policy samples $\{\theta_{i}\}$
29
30  Update critics $Q_{1},Q_{2}$ using averaged targets
31
 If: adaptation steps completed then $\text{actor\_frozen}\leftarrow\text{False}$
32
 /* PAC-Bayes Update Cycle */
33 if $t\bmod\text{pb\_update\_freq}=0$ then
    $\mathfrak{D}_{\text{rollouts}}\leftarrow\text{collect\_fresh\_rollouts}()$
    // With the current policy
34    $\tau_{\min}\leftarrow\text{estimate\_mixing\_time}(\mathfrak{D}_{\text{rollouts}})$
35
36    $\mathfrak{D}_{\text{train}},\mathfrak{D}_{\text{test}}\leftarrow\text{split}(\mathfrak{D}_{\text{rollouts}})$
37
38     Compute discounted returns $G_{\text{IS}}$ with importance sampling on $\mathfrak{D}_{\text{train}}$
39
    /* Alternating optimization */
40    for epoch $1,\ldots,\text{pb\_epochs}$ do
       /* Optimize posterior parameters for fixed $\lambda$ */
       $\sigma,\upsilon\leftarrow\arg\min_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\bm{(\sigma,\upsilon)}}}\mathcal{J}(\rho,\lambda)$
       // inequality 14
41
       /* Optimize $\lambda$ for fixed posterior */
       $\lambda\leftarrow\arg\min_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\bm{\lambda^{\prime}}}}\mathcal{J}(\rho,\lambda^{\prime})$
       // inequality 14
42
43      end for
44
45    $\text{bound}\leftarrow\text{compute\_pac\_bayes\_bound}(\mathfrak{D}_{\text{test}},\text{KL}(\rho\|\mu),\tau_{\min})$
46
    $\text{load\_policy\_params}(\upsilon)$
    // Sync actor to posterior mean
47
    $\text{actor\_frozen}\leftarrow\text{True}$
    // Initiate critic adaptation
48
49   end if
 /* Prior Reset for Maintained Exploration */
50 if $t\bmod\text{pb\_reset\_freq}=0$ then
    $\upsilon_{0}\leftarrow\iota\cdot\upsilon+(1-\iota)\cdot\upsilon_{0};\quad\sigma_{0}\leftarrow\iota\cdot\sigma+(1-\iota)\cdot\sigma_{0}$
    // Moving average prior update
51
52     Linearly decay $\iota$ and $\epsilon_{\text{explore}}$
53   end if
54
55 end for
Algorithm 1 PAC-Bayes Soft Actor-Critic (PB-SAC)
Appendix G More Results
This appendix presents extended experimental results that complement the main paper findings. Figure 4 shows detailed performance comparisons on Hopper-v5 and Walker2d-v5 environments, demonstrating both the algorithm comparisons (panel a) and the PAC-Bayes analysis (panel b). The empirical discounted return tracks closely with our certified lower bound, validating our analysis above.
(a) Algorithm comparisons
(b) PAC-Bayes analysis
Figure 4: (a) Performance comparison between our   PB-SAC, its baseline   SAC, and   PBAC from Tasdighi et al. (2025); (b) PAC-Bayes analysis of PB-SAC across environments. The empirical discounted return (dashed line) corresponds to $\mathbb{E}_{\theta\sim\rho}[-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]$, and the certified discounted return (solid line) corresponds to the lower bound on $\mathbb{E}_{\theta\sim\rho}[-\mathcal{L}(\theta)]$ provided by Theorem 3.2 (after rearranging the terms).
G.1 Sparse-reward setting
(a) Algorithm comparisons
(b) PAC-Bayes analysis
Figure 5: (a) Performance comparison between our   PB-SAC, its baseline   SAC, and   PBAC from Tasdighi et al. (2025); (b) PAC-Bayes analysis of PB-SAC across environments. The empirical discounted return (dashed line) corresponds to $\mathbb{E}_{\theta\sim\rho}[-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]$, and the certified discounted return (solid line) corresponds to the lower bound on $\mathbb{E}_{\theta\sim\rho}[-\mathcal{L}(\theta)]$ provided by Theorem 3.2 (after rearranging the terms).
(a) Mixing time: 1
(b) Mixing time: 50
(c) Mixing time: 100
(d) Mixing time: 300
(e) Mixing time: 500
(f) Mixing time: 1000
Figure 6: PAC-Bayesian bound behaviour with different fixed mixing time estimates on HalfCheetah-v5. Empirical discounted return (dashed) and certified discounted return (solid) are shown for mixing times ranging from 1 (a) to 1000 (f). Smaller mixing times yield tighter bounds but risk overconfidence if underestimated, while larger values provide conservative but still meaningful certificates. The algorithm maintains practical utility even with substantial mixing time overestimation, demonstrating robustness of the framework.
Appendix H Ablation study
(a) Adaptive Sampling
(b) Posterior-Guided Exploration
Figure 7: Ablations
Figure 8: KL divergence evolution
Appendix I Hyperparameter Selection
We carefully selected hyperparameters for our PAC-Bayes Soft Actor-Critic (PB-SAC) implementation to balance performance, sample efficiency, and theoretical guarantees. the common parameters with SAC are left unchanged, while we take the original hyperparameters of PBAC from the paper of Tasdighi et al. (2025). Table 1 summarizes it all.
Table 1: Hyperparameter Comparison for MuJoCo Continuous Control Tasks
Hyperparameter PB-SAC (Our Algorithm) SAC (Baseline) PBAC
Common SAC Parameters
Total Timesteps $\times 10^{6} 1×1$$\times 10^{6}$ $1\times 10^{6}$
Discount Factor ($\gamma$) $0.99$ $0.99$ $0.99$
Soft Update Coefficient ($\tau$) $0.005$ $0.005$ $0.005$
Batch Size $256$ $256$ $256$
Replay Buffer Size $\times 10^{6} 1×1$$\times 10^{6}$ $1\times 10^{6}$
Initial Temperature ($\alpha$) $0.2$ $0.2$ $0.2$
Temperature Learning Rate $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
Target Update Frequency 11 11 11
Algorithm-Specific Parameters
Actor Learning Rate $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
Critic Learning Rate $\times 10^{-3} 1×1$$\times 10^{-3}$ $3\times 10^{-4}$
Learning Starts 5,0005{,}000 5,0005{,}000 10,00010{,}000
Training Frequency 22 22 11
Automatic $\alpha$ Tuning $\times$
Multi-Head Architecture $\times$ $\times$
Ensemble of Critics $\times$ $\times$ $(10)$
Number of Heads 11 11 $10$
Network Architecture
Policy Hidden Layers $[256,256]$ $[256,256]$ $[256,256]$
Q-Function Hidden Layers $[256,256]$ $[256,256]$ $[256,256]$
Activation Function ReLU ReLU CReLU
PAC-Bayes Specific (PB-SAC Only)
Failure Probability ($\delta$) $0.1$ - -
Initial Std Dev $0.01$ - -
PB Update Frequency 20,00020{,}000 - -
Actor Freeze Frequency 2020 - -
Adaptation samples $256$ - -
PB Rollout Trajectories $100$ - -
PB Rollout Steps per Trajectory $500$ - -
PBAC Specific
Bootstrap Rate - - $0.05$
Posterior Sampling Rate - - 55
Prior Scaling - - $5.0$